OpenStudy (waheguru):

need some help im getting 311

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

???

OpenStudy (karatechopper):

311?

OpenStudy (waheguru):

im posting the question sorry

OpenStudy (waheguru):

here we are

OpenStudy (anonymous):

Do you have the formula for the surface area of a cylinder?

OpenStudy (waheguru):

2pirsq plus 2pirh

OpenStudy (anonymous):

Surface area of fully cylinder = circumference * length. Here, you have half of it. So, 1/2*circumference*length = 1/2 * pi * 6 * 30 = 90 * pi

jhonyy9 (jhonyy9):

a.

OpenStudy (anonymous):

waheguru - that is when including bottom and top circles. In this case, the roof has no such thing. So, you just need 1/2 circumference * length.

OpenStudy (asnaseer):

@GT is right

OpenStudy (waheguru):

oh so i included the bottom???

OpenStudy (anonymous):

Yeah. Don't.

OpenStudy (waheguru):

k so this is what i did 2 x pi x 3^2 + 2 x pi x 3 x 30 and divide by 2

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

Remove 3*pi*r^2 and the rest is good.

OpenStudy (waheguru):

y

jhonyy9 (jhonyy9):

a. is right sure

OpenStudy (anonymous):

Because the roof does not have a base and top (according to the problem description).

OpenStudy (asnaseer):

the question is asking for the area of the roofs curved surface - that does not include the two ends of the greenhouse

OpenStudy (waheguru):

so just the lateral

OpenStudy (anonymous):

yes.

OpenStudy (waheguru):

thx brozz

OpenStudy (asnaseer):

the curved surface of a cylinder has a surface area of \(2\pi rh\). in this case you have half a cylinder, so the surface area becomes \(\pi rh\)

OpenStudy (anonymous):

HI GUYS

OpenStudy (waheguru):

just practising for grade 9 EQAO

OpenStudy (asnaseer):

what is EQAO?

OpenStudy (waheguru):

evil questions attacking ontario lol im kidding its like an exam for grade 9 students

OpenStudy (asnaseer):

lol! :-D - good luck with them...

OpenStudy (waheguru):

thx

jhonyy9 (jhonyy9):

is half perimeter of base multiply lenght=surface area

jhonyy9 (jhonyy9):

this very easy

OpenStudy (asnaseer):

@jhonyy9 - things are always easy when you know how to do them - try and recall the days when you were also at the lower end of the learning curve :-)

OpenStudy (asnaseer):

like we ALL were

jhonyy9 (jhonyy9):

yes right but this is the same like a rectangle curved on the base